3. If the return on assets is .15 and average assets are 100, the EPS is:

  1. 15

  2. 115

  3. 85

  4. Unable to tell from data given

7. Arnold llc has base sales of 100 and rent expense is 12% of sales. Assuming a 15% increase in sales, rent expense for the first pro forma year is:

  1. 13.80

  2. 14.80

  3. 15.80

  4. 16.80
Answer #1

(3) Return on Assets = Net Income / Average Assets

=> Net Income = Return on Assets * Average Assets = 0.15 * 100 = 15

EPS = (Net Income - Dividends) / Outstanding Shares

However, we do not have information on Dividends and Number of Outstanding shares.

Hence (d) Unable to tell from data given

(4) Sales = 100

Rent Expense = 12% of Sales = 12

New Sales = 100 + 0.15*100 = 115

New Rent Expense = 12% of New Sales = 0.12 *115 = 13.80

Hence, correct option is (a) 13.80